# Almost Surely Asymptotically Constant Graph Neural Networks

### v2

@@ Abstract
We present a new angle on the expressive power of graph neural networks (GNNs) by studying how the predictions of a GNN probabilistic classifier evolve as we apply it on larger graphs drawn from some random graph model. We show that the output converges to a constant function, which upper-bounds what these classifiers can uniformly express. This strong convergence phenomenon applies to a very wide class of GNNs, including state of the art models, with aggregates including mean and the attention-based mechanism of graph transformers. Our results apply to a broad class of random graph models, including sparse and dense variants of the Erdős–Rényi model, the stochastic block model, and the Barabási-Albert model. We empirically validate these findings, observing that the convergence phenomenon appears not only on random graphs but also on some real-world graphs.
@@ 1 Introduction
Graph neural networks (GNNs) [47, 21] have become prominent for graph ML with applications in domains such as life sciences [50, 11, 25, 58]. Their empirical success motivated work investigating their theoretical properties, pertaining to their expressive power [55, 42, 4, 8, 1, 46, 23, 19, 57], generalization capabilities [18, 33, 40, 48, 43], and convergence properties [28, 27, 40, 32, 2].
We consider GNNs acting as probabilistic graph classifiers and ask: how do the predictions of GNNs acting as probabilistic graph classifiers evolve as we apply them on larger graphs drawn from a random graph model?
Our study gives a surprising answer to this question: the output of many probabilistic classifiers eventually become independent of their inputs – each model eventually outputs the same prediction probabilities on all graphs – as graph sizes increase (Figure 1). This “almost sure convergence” to a constant distribution is much stronger than convergence to some limit object [28, 27, 32]. The immediate consequence of this strong convergence phenomenon is to upper bound the uniform expressiveness of the considered model architectures: these architectures can uniformly express only the probabilistic classifiers that are almost surely asymptotically constant.
Scope of the result. The core approach in graph ML is based on iteratively updating node representations of an input graph by an aggregate of messages flowing from the node’s neighbours [20]. This approach can be extended to global aggregates [5]. Our main result holds for all architectures that use weighted mean as an aggregation function and it is extremely robust in the following two dimensions:
@@
    Model architectures: Our result is very general and abstracts away from low-level architectural design choices. To achieve this, we introduce an aggregate term language using weighted mean aggregation and provide an “almost sure optimization” result for this language: our result states that every term in the language can be simplified to Lipschitz functions for most inputs. Thus, any architecture that can be expressed in this language follows the same convergence law. This includes graph attention networks (GATs) [52], as well as popular (graph) transformers, such as the General, Powerful, Scalable Graph Transformer (GPS) with random walk encodings [44]. The term language can seemlessly capture common design choices, such as skip and jumping knowledge connections [54], or global aggregation schemes [5].
@@
  • We show that a diverse class of architectures acting as probabilistic graph classifiers can be expressed in the term language (Section 4) and derive “almost sure convergence” as corollaries to our main result (Corollary 5.2). The results are robust to many practical architectural design choices and even hold for architectures using mixtures of layers from different architectures.
  • We validate these results empirically, showing the convergence of these probabilistic graph classifiers in practice (Section 6) on graphs drawn from the random models studied. In addition, we probe the real-world import of our results by testing for convergence on a dataset with varying size dataset splits. Across all experiments we observe rapid convergence to a constant distribution. Interestingly, we note some distinctions between the convergence of the sparse and non-sparse Erdős-Rényi model, which we can relate to the proof strategies for our convergence laws.
@@ 2 Related work
Uniform expressiveness. The expressive power of MPNNs is studied from different angles, including their power in terms of graph distinguishability [55, 42]. The seminal results of Xu et al. [55], Morris et al. [42] show that MPNNs are upper bounded by the 1-dimensional Weisfeiler Leman graph isomorphism test (1-WL) in terms of graph distinguishability. WL-style expressiveness results are inherently non-uniform, i.e., the model construction is dependent on the graph size. There are also recent studies that focus on uniform expressiveness [4, 45, 2]. In particular, Adam-Day et al. [2] investigate the uniform expressive power of GNNs with randomized node features, which are known to be more expressive in the non-uniform setting [1, 46]. They show that for classical Erdős-Rényi graphs, GNN binary classifiers display a zero-one law, assuming certain restrictions on GNN weights and the random graph model. We focus on probabilistic classifiers, where their results do not apply, while dealing with a wider class of random graph models, subsuming popular architectures such as graph transformers.
@@ 3.2 Graph neural networks and graph transformers
$\small\textsc{Att}(v)=\sum_{u\in V_{G}}\frac{{\exp\left(\mathrm{scale}\left({\bm{h}}_{v}^{(\ell)},{\bm{h}}_{u}^{(\ell)}\right)\right)}}{\sum\limits_{w\in V_{G}}\exp\left(\mathrm{scale}\left({\bm{h}}_{v}^{(\ell)},{\bm{h}}_{w}^{(\ell)}\right)\right)}{\bm{W}}^{(\ell)}_{V}{\bm{h}}_{u}^{(\ell)}$
@@
Probabilistic classifiers. All of these models can be used as probabilistic graph classifiers. We only need to ensure that the final layer is a softmax or sigmoid applied to pooled representations.
@@ 4 Model architectures via term languages
  1. The basic terms are of the form $\mathrm{H}(x)$, representing the features of the node $x$, and constants ${\bm{c}}$.
  2. Let $h\colon\mathbb{R}^{d}\to(0,\infty)^{d}$ be a function which is Lipschitz continuous on every compact domain. Given terms $\tau$ and $\pi$, the local $h$-weighted mean for node variable $x$ is:
@@
  3. Terms are closed under applying a function symbol for each Lipschitz continuous $F\colon\mathbb{R}^{d\times k}\to\mathbb{R}^{d}$.
The weighted mean operator takes an average of the values returned by $\pi$, using $\tau$ to perform a weighting and normalizing the values of $\pi$ using $h$ to ensure that we are not dividing by zero.
@@
We make the interpretation of the terms precise as follows.
@@
  1. $\llbracket{\bm{c}}(\bar{u})\rrbracket_{G}={\bm{c}}$ for any constant ${\bm{c}}$.
  2. $\llbracket\mathrm{H}(x_{i})(\bar{u})\rrbracket_{G}=H_{G}(u_{i})$ for the $i$th node’s features.
  3. $\llbracket F(\tau_{1},\ldots,\tau_{k})(\bar{u})\rrbracket_{G}=F(\llbracket\tau_{1}(\bar{u})\rrbracket_{G},\ldots,\llbracket\tau_{k}(\bar{u})\rrbracket_{G})$ for any function symbol $F$.
  4. For any term composed using $\star$:
    $\left\llbracket\sum_{y\in\mathcal{N}(x_{i})}\tau\star h(\pi)(\bar{u})\right\rrbracket_{G}=\left\{\begin{array}[]{cc}\nicefrac{{\sum\limits_{v\in\mathcal{N}(u_{i})}\llbracket\tau(\bar{u},v)\rrbracket_{G}h(\llbracket\pi(\bar{u},v)\rrbracket_{G})}}{{\sum\limits_{v\in\mathcal{N}(u_{i})}h(\llbracket\pi(\bar{u},v)\rrbracket_{G})}}&\mbox{if $\operatorname{\mathcal{N}}(u_{i})\neq\emptyset$};\\ {\bf 0}&\mbox{if $\operatorname{\mathcal{N}}(u_{i})=\emptyset$}.\end{array}\right.$
@@ 4. For any term composed using $\star$:
$\llbracket\text{rw}(x_{i})(\bar{u})\rrbracket=[\text{rw}_{u_{i},1},\ldots,\text{rw}_{u_{i},d}]$
@@ 4.1 How powerful is the term language?
  1. Let $\tau^{(0)}(x)\coloneqq\mathrm{H}(x)$. (Initialization)
  2. For $1\leq\ell<L$, define $\tau^{(\ell+1)}(x)\coloneqq\small\textsc{Upd}^{(\ell)}\left(x,\small\textsc{Mean}_{y\in\mathcal{N}(x)}\tau^{(\ell-1)}(y)\right)$. (Layers)
  3. The final graph representation is computed as $\tau\coloneqq\small\textsc{Mean}_{x}\tau^{(L)}(x)$. (Mean pooling)
Using this construction, and letting ${\bm{h}}_{v}^{(\ell)}$ be the representation of the node $v$ in a graph $G$ computed in layer $\ell$ of the MPNN, it is straightforward to verify inductively that $\llbracket\tau^{(\ell)}(v)\rrbracket_{G}={\bm{h}}_{v}^{(\ell)}$.
@@ 5 Convergence theorems
Concretely, this result shows that for any probabilistic classifier which can be expressed within the term language, when drawing graphs from any of these distributions, eventually the output of the classifier will be the same regardless of input graph, asymptotically almost surely. Thus, the only classifiers which can be expressed by such models are those which are asymptotically constant.
@@ 5.1 Overview of the technical constructions for the denser cases
In the denser cases, we first show that we have stable behaviour in the node degrees. The precise formalization will be slightly different in each case, but in each one neighbourhoods grow at least logarithmically with graph size:
Lemma 5.3 (Log neighbourhood bound).
@@
The key lemma below shows that to each term $\pi$ we can associate a feature-type controller $e_{\pi}$ which captures the asymptotic behaviour of $\pi$, in the sense that with high probability, for most of the tuples $\bar{u}$ the value of $e_{\pi}(H_{G}(\bar{u}),\mathrm{GrTp}(\bar{u}))$ is close to $\llbracket\pi(\bar{u})\rrbracket$.
Lemma 5.4 (Key Lemma).
@@ 5.2 Overview of the technical constructions for the sparser cases
In the sparser cases, we proceed a bit differently. Instead of graph types over $\bar{u}$, which only specify graph relations among the $\bar{u}$, we require descriptions of local neighbourhoods of $\bar{u}$.
@@
Let $G$ be a graph, $\bar{u}$ a tuple of nodes in $G$ and $\ell\in\mathbb{N}$. The $\ell$-neighbourhood of $\bar{u}$ in $G$, denoted $\operatorname{\mathcal{N}}_{\ell}(\bar{u})$ is the subgraph of $G$ induced by the nodes which have a path of length at most $\ell$ to some node in $\bar{u}$.
The “types” are now graphs $T$ with $k$ distinguished elements $\bar{w}$, which we call $k$-rooted graphs. The combinatorial tool here is a fact known in the literature as ‘weak local convergence’: the percentage of local neighbourhoods of any given type converges.
@@
Our inductive invariant, analogous to Lemma 5.4 states that every term can be approximated by a Lipschitz function of the features and the neighbourhood type.
Lemma 5.7.
@@ 6 Experimental evaluation
We first empirically verify our findings on random graphs and then on a real-world graph to answer the following questions: Q1. Do we empirically observe convergence? Q2. What is the impact of the different weighted mean aggregations on the convergence? Q3. What is the impact of the graph distribution on the convergence? Q4. Can this phenomena arise within large real-world graphs?
@@
  1. We consider five models with the same architecture, each having randomly initialized weights, utilizing a $\operatorname{ReLU}$ non-linearity, and applying a softmax function to their outputs. Each model uses a hidden dimension of $128$, $3$ layers and an output dimension of $5$.
  2. We experiment with distributions $\mathrm{ER}(n,p(n)=0.1)$, $\mathrm{ER}(n,p(n)=\frac{\log{n}}{n})$, $\mathrm{ER}(n,p(n)=\frac{1}{50n})$, $\mathrm{BA}(n,m=5)$. We also experiment with an SBM of 10 communities with equal size, where an edge between nodes within the same community is included with probability $0.7$ and an edge between nodes of different communities is included with probability $0.1$.
To understand the behaviour of the respective models, we will draw larger and larger graphs from the graph distributions. We use five different models to ensure this is not a model-specific behaviour.

### v3

@@ Abstract
We present a new angle on the expressive power of graph neural networks (GNNs) by studying how the predictions of real-valued GNN classifiers, such as those classifying graphs probabilistically, evolve as we apply them on larger graphs drawn from some random graph model. We show that the output converges to a constant function, which upper-bounds what these classifiers can uniformly express. This strong convergence phenomenon applies to a very wide class of GNNs, including state of the art models, with aggregates including mean and the attention-based mechanism of graph transformers. Our results apply to a broad class of random graph models, including sparse and dense variants of the Erdős–Rényi model, the stochastic block model, and the Barabási-Albert model. We empirically validate these findings, observing that the convergence phenomenon appears not only on random graphs but also on some real-world graphs.
@@ 1 Introduction
Graph neural networks (GNNs) [44, 21] have become prominent for graph machine learning with applications in domains such as life sciences [47, 11, 25, 55]. Their empirical success motivated work investigating their theoretical properties, pertaining to their expressive power [52, 39, 4, 8, 1, 43, 23, 19, 54], generalization capabilities [18, 33, 38, 45, 40], and convergence properties [28, 27, 38, 32, 2].
We consider GNNs outputting real-valued vectors, such as those which classify graphs probabilistically, and ask: how do the outputs of these GNNs evolve as we apply them on larger graphs drawn from a random graph model?
Our study provides a surprising answer to this question: the output of many GNNs eventually become independent of their inputs – each model eventually outputs the same values on all graphs – as graph sizes increase (Figure 1). This “almost sure convergence” to a constant distribution is much stronger than convergence to some limit object [28, 27, 32]. The immediate consequence of this strong convergence phenomenon is to upper bound the uniform expressiveness of the considered model architectures: these architectures can uniformly express only the classifiers that are almost surely asymptotically constant. In other words, our results provide impossibility results for what tasks are in principle learnable by GNNs. While our top-level results are for graph classification, in the process we provide strong limitations on what node- and edge-classification can be performed by GNNs on random graphs: see, for example Theorem 5.3.
Scope of the result. The core approach in graph machine learning is based on iteratively updating node representations of an input graph by an aggregate of messages flowing from the node’s neighbours [20]. This approach can be extended to global aggregates [5]. Our main result holds for all architectures that use weighted mean as an aggregation function and it is extremely robust in the following two dimensions:
@@
    Model architectures: Our result is very general and abstracts away from low-level architectural design choices. To achieve this, we introduce an aggregate term language using weighted mean aggregation and provide an “almost sure optimization” result for this language: our result states that every term in the language can be simplified to Lipschitz functions for most inputs. Thus, any architecture that can be expressed in this language follows the same convergence law. This includes graph attention networks (GATs) [49], as well as popular (graph) transformers, such as the General, Powerful, Scalable Graph Transformer (GPS) with random walk encodings [41]. The term language can seamlessly capture common design choices, such as skip and jumping knowledge connections [51], or global aggregation schemes [5].
@@
  • We show that a diverse class of architectures acting as real-valued graph classifiers can be expressed in the term language (Section 4). In Section 5 we present “almost sure convergence” results for our term language, from which we derive results about GNNs (Corollary 5.2). The results are robust to many practical architectural design choices and even hold for architectures using mixtures of layers from different architectures. We also show strong convergence results for real-valued node classifiers in many graph models.
  • We validate these results empirically, showing the convergence of these graph classifiers in practice (Section 6) on graphs drawn from the random models studied. In addition, we probe the real-world significance of our results by testing for convergence on a dataset with varying size dataset splits. Across all experiments we observe rapid convergence to a constant distribution. Interestingly, we note some distinctions between the convergence of the sparse and non-sparse Erdős-Rényi model, which we can relate to the proof strategies for our convergence laws.
@@ 2 Related work
Uniform expressiveness. The expressive power of MPNNs is studied from different angles, including their power in terms of graph distinguishability [52, 39]. The seminal results of Xu et al. [52], Morris et al. [39] show that MPNNs are upper bounded by the 1-dimensional Weisfeiler Leman graph isomorphism test (1-WL) in terms of graph distinguishability. WL-style expressiveness results are inherently non-uniform, i.e., the model construction is dependent on the graph size. There are also recent studies that focus on uniform expressiveness [4, 42, 2]. In particular, Adam-Day et al. [2] investigate the uniform expressive power of GNNs with randomized node features, which are known to be more expressive in the non-uniform setting [1, 43]. They show that for classical Erdős-Rényi graphs, GNN binary classifiers display a zero-one law, assuming certain restrictions on GNN weights and the random graph model. We focus on real-valued classifiers, where their results do not apply, while dealing with a wider class of random graph models, subsuming popular architectures such as graph transformers.
@@ 3.2 Graph neural networks and graph transformers
$\small\textsc{Att}(v)=\sum_{u\in V_{G}}\frac{{\exp\left(\mathrm{scale}\left({\bm{h}}_{v}^{(\ell)},{\bm{h}}_{u}^{(\ell)}\right)\right)}}{\sum\limits_{w\in V_{G}}\exp\left(\mathrm{scale}\left({\bm{h}}_{v}^{(\ell)},{\bm{h}}_{w}^{(\ell)}\right)\right)}{\bm{W}}^{(\ell)}{\bm{h}}_{u}^{(\ell)}$
@@
Probabilistic classifiers. We are looking at models that produce a vector of reals on each graph. All of these models can be used as probabilistic graph classifiers. We only need to ensure that the final layer is a softmax or sigmoid applied to pooled representations.
@@ 4 Model architectures via term languages
  • The basic terms are of the form $\mathrm{H}(x)$, representing the features of the node $x$, and constants ${\bm{c}}$.
  • Let $h\colon\mathbb{R}^{d}\to(0,\infty)^{d}$ be a function which is Lipschitz continuous on every compact domain. Given terms $\tau$ and $\pi$, the local $h$-weighted mean for node variable $x$ is:
@@
  • Terms are closed under applying a function symbol for each Lipschitz continuous $F\colon\mathbb{R}^{d\times k}\to\mathbb{R}^{d}$ for any $k\in\mathbb{N}^{+}$.
The weighted mean operator takes a weighted average of the values returned by $\tau$. It uses $\pi$ to perform a weighting, normalizing the values of $\pi$ using $h$ to ensure that we are not dividing by zero (see below for the precise definition).
@@
We make the interpretation of the terms precise as follows. See Figure 2 for a graphical example of evaluating a term on a graph.
@@
  • $\llbracket{\bm{c}}(\bar{u})\rrbracket_{G}={\bm{c}}$ for any constant ${\bm{c}}$.
  • $\llbracket\mathrm{H}(x_{i})(\bar{u})\rrbracket_{G}=H_{G}(u_{i})$ for the $i$th node’s features.
  • $\llbracket F(\tau_{1},\ldots,\tau_{k})(\bar{u})\rrbracket_{G}=F(\llbracket\tau_{1}(\bar{u})\rrbracket_{G},\ldots,\llbracket\tau_{k}(\bar{u})\rrbracket_{G})$ for any function symbol $F$.
  • For any term composed using $\star$:
    $\left\llbracket\sum_{y\in\mathcal{N}(x_{i})}\tau\star h(\pi)(\bar{u})\right\rrbracket_{G}=\left\{\begin{array}[]{cc}\frac{\sum\limits_{v\in\mathcal{N}(u_{i})}\llbracket\tau(\bar{u},v)\rrbracket_{G}h(\llbracket\pi(\bar{u},v)\rrbracket_{G})}{\sum\limits_{v\in\mathcal{N}(u_{i})}h(\llbracket\pi(\bar{u},v)\rrbracket_{G})}&\mbox{if $\operatorname{\mathcal{N}}(u_{i})\neq\emptyset$};\\ {\bf 0}&\mbox{if $\operatorname{\mathcal{N}}(u_{i})=\emptyset$}.\end{array}\right.$
@@
Figure 2: Evaluation of the term $\sum_{x\in\mathcal{N}(y)}(2\mathrm{H}(x)+2)\star 1.0$ on a small graph with scalar features. The term computes the mean of $z\mapsto 2z+2$ on each of a node’s neighbours. As each sub-term has one free variable, we can represent the intermediate results as scalar values for each node.
@@
$\llbracket\text{rw}(x_{i})(\bar{u})\rrbracket_{G}=[\text{rw}_{u_{i},1},\ldots,\text{rw}_{u_{i},d}]$
@@ 4.1 How powerful is the term language?
  • Initialization. Let $\tau^{(0)}(x)\coloneqq\mathrm{H}(x)$. Then the value $\llbracket\tau^{(0)}(x)(u)\rrbracket_{G}$ at a node $u$ is the initial node representation $H(u)$.
  • Layers. For $1\leq\ell<L$, define $\tau^{(\ell+1)}(x)\coloneqq\small\textsc{Upd}^{(\ell)}\left(x,\small\textsc{Mean}_{y\in\mathcal{N}(x)}\tau^{(\ell)}(y)\right)$. Then the value at node $u$ is the following, which conforms with the inductive construction of the MPNN:
    $\llbracket\tau^{(\ell+1)}(x)(u)\rrbracket_{G}=\small\textsc{Upd}^{(\ell)}\left(\llbracket\tau^{(\ell)}(x)(u)\rrbracket_{G},\frac{1}{|\mathcal{N}(u)|}\sum_{v\in\mathcal{N}(u)}\llbracket\tau^{(\ell)}(x)(v)\rrbracket_{G}\right)$
  • Final mean pooling. The final graph representation is computed as $\tau\coloneqq\small\textsc{Mean}_{x}\tau^{(L)}(x)$.
@@ 5 Convergence theorems
Concretely, this result shows that for any probabilistic classifier, or other real-valued classifier, which can be expressed within the term language, when drawing graphs from any of these distributions, eventually the output of the classifier will be the same regardless of the input graph, asymptotically almost surely. Thus, the only probabilistic classifiers which can be expressed by such models are those which are asymptotically constant.
@@ 5.1 Overview of the technical constructions for the denser cases
The key theorem below shows that to each term $\pi$ we can associate a feature-type controller $e_{\pi}$ which captures the asymptotic behaviour of $\pi$, in the sense that with high probability, for most of the tuples $\bar{u}$ the value of $e_{\pi}(H_{G}(\bar{u}),\mathrm{GrTp}(\bar{u}))$ is close to $\llbracket\pi(\bar{u})\rrbracket$.
Theorem 5.3 (Aggregate Elimination for Non-Sparse Graphs).
@@
Theorem 5.3 highlights an advantage of working with a term language having nice closure properties, rather than directly with GNNs: it allows us to use induction on term construction, which may be more natural and more powerful than induction on layers. Theorem 5.3 also gives strong limitations on node and link classification using GNNs: on most nodes in (non-sparse) random graphs, GNNs can only classify based on the features of a node, they cannot make use of any graph structure.
@@ 5.2 Overview of the technical constructions for the sparser cases
In the sparser cases, the analysis is a bit more involved. Instead of graph types over $\bar{u}$, which only specify graph relations among the $\bar{u}$, we require descriptions of local neighbourhoods of $\bar{u}$.
@@
Let $G$ be a graph, $\bar{u}$ a tuple of nodes in $G$ and $\ell\in\mathbb{N}$. The $\ell$-neighbourhood of $\bar{u}$ in $G$, denoted $\operatorname{\mathcal{N}}_{\ell}(\bar{u})$ is the subgraph of $G$ induced by the nodes of distance at most $\ell$ from some node in $\bar{u}$.
The “types” are now graphs $T$ with $k$ distinguished elements $\bar{w}$, which we call $k$-rooted graphs. Two $k$-rooted graphs $(T,\bar{w})$ and $(U,\bar{z})$ are isomorphic is there is a structure-preserving bijection $T\to U$ which maps $\bar{w}$ to $\bar{z}$. The combinatorial tool here is a fact known in the literature as ‘weak local convergence’: the percentage of local neighbourhoods of any given type converges.
@@
Our “aggregate elimination”, analogous to Theorem 5.3, states that every term can be approximated by a Lipschitz function of the features and the neighbourhood type.
Theorem 5.6 (Aggregate Elimination for Sparser Graphs).
@@
Compared to Theorem 5.3 the result is much less limiting in what node-classifying GNNs can express. Although combining the sparse and non-sparse conditions covers many possible growth rates, it is not true that one gets convergence for Erdős-Rényi with arbitrary growth functions:
Theorem 5.7.
There are functions $p(n)$ converging to zero and a term $\tau$ in our language such that $\tau$ does not converge even in distribution (and hence does not converge a.a.s.) over ER random graphs with growth rate $p$.
Proof.
Let $p$ alternate between $\frac{1}{2}$ on even $n$ and $\frac{1}{n}$ on odd $n$. Consider $\tau_{1}(x)$ that returns 00 if $x$ has a neighbour and $1$ otherwise, and let $\tau$ be the global average of $\tau_{1}$. So $\tau$ is the percentage of isolated nodes in the graph. Then $\tau$ clearly goes to zero in the non-sparse case. However the probability that a particular node is not isolated in a random sparse graph of size $n$ is $(1-\frac{1}{n})^{n-1}$, which goes to $\frac{1}{e}<1$. Thus $\llbracket\tau\rrbracket$ diverges on $\mathrm{ER}(n,p(n))$. ∎
@@ 6 Experimental evaluation
We first empirically verify our findings on random graphs and then on a real-world graph to answer the following questions: Q1. Do we empirically observe convergence? Q2. What is the impact of the different weighted mean aggregations on the convergence? Q3. What is the impact of the graph distribution on the convergence? Q4. Can these phenomena arise within large real-world graphs?
@@
  • We consider five models with the same architecture, each having randomly initialized weights, utilizing a $\operatorname{ReLU}$ non-linearity, and applying a softmax function to their outputs. Each model uses a hidden dimension of $128$, $3$ layers and an output dimension of $5$.
  • We experiment with distributions $\mathrm{ER}(n,p(n)=0.1)$, $\mathrm{ER}(n,p(n)=\frac{\log{n}}{n})$, $\mathrm{ER}(n,p(n)=\frac{1}{50n})$, $\mathrm{BA}(n,m=5)$. We also experiment with an SBM of 10 communities with equal size, where an edge between nodes within the same community is included with probability $0.7$ and an edge between nodes of different communities is included with probability $0.1$.
  • We draw graphs of sizes up to 10,000, where we take 100 samples of each graph size. Node features are independently drawn from $U[0,1]$ and the initial feature dimension is $128$.
To understand the behaviour of the respective models, we will draw larger and larger graphs from the graph distributions. We use five different models to ensure this is not a model-specific behaviour. Further experimental results are reported in Appendix F.